OpenStudy (anonymous):

Let f(x) = 9x - 2 and g(x) = -x + 3. Find f(g(x)).

OpenStudy (anonymous):

-9x - 2 -9x + 5 -9x + 25 -9x + 27

OpenStudy (anonymous):

IS IT A ?

OpenStudy (anonymous):

@Faceomo @eric_d

OpenStudy (eric_d):

@kirbykirby is here

OpenStudy (anonymous):

OH I DIDNT KNOW SORRY

OpenStudy (kirbykirby):

\(f(\color{red}{x})=9\color{red}{x}-2\) \(g(x)=-x+3\) \(f(\color{red}{g(x)})= 9(\color{red}{g(x)})-2=9(\color{red}{-x+3})-2\)

OpenStudy (kirbykirby):

Do you see how that works? Whatever is in between f ( ...) , plug it where you see it in the function. In f(x) = 9x + 2... whatever is in ( ) before f, you substitute in your function. so if you have f(g(x)), you substitute wherever you see "x" with "g(x)"

OpenStudy (anonymous):

KIRBY DO I DO THE ALGEBRA STILL ?

OpenStudy (kirbykirby):

yes, you must simplify the answer in order to get one of the choices you are given

OpenStudy (anonymous):

THANKS U HAVE BEEN A LOT OF HELP :}

OpenStudy (kirbykirby):

your welcome =]

OpenStudy (anonymous):

@fateal @Alfa_female

OpenStudy (kirbykirby):

No, if you plug in f into the g, that would be \(g(f(x))\)

OpenStudy (anonymous):

KIRBY I STILL DONT GET IT

OpenStudy (anonymous):

AND ITS PISSING ME OFF

OpenStudy (kirbykirby):

Did you understand how I got this? \[ f(\color{red}{g(x)})= 9(\color{red}{g(x)})-2=9(\color{red}{-x+3})-2\]

OpenStudy (anonymous):

I THINK SO

OpenStudy (anonymous):

ok i honestly thought i could help you buti cant i have done these as much and frankly not to this extent sorry @taylor12344

OpenStudy (anonymous):

haven't

OpenStudy (fateal):

i can help but this computer is killing me :/ slow as f >.<

OpenStudy (anonymous):

OK

OpenStudy (kirbykirby):